# P.O.Box 19395-5531, Tehran, Iran

### v1

@@ Abstract
Despite previous results that rule out the possibility of a static black hole horizon in cosmology we present a black hole metric that has a static horizon whilst mimicking the cosmological behavior at large scales away from the black hole horizon. By a suitable choice of coordinates, we show that it is possible to have a static black hole horizon in cosmology that does not suffer from issues like singularities in curvature invariants at the horizon. The resulting metric is consistent with the Schwarzschild-de Sitter metric for constant Hubble parameter. It should be noted that this metric does not lead to isotropic pressure, i.e. $p_{\perp}\neq p_{r}$ for radial distances comparable to Schwarzschild radius, but for large $r$ the stress tensor (assuming general relativity) tends to the cosmological value. Turning to a more realistic framework, we study the cases of black hole$+$cosmological matter and black hole$+$matter$+$cosmological constant where we assume the pressure component takes its cosmological value and show that although the resulting metric does not have a static horizon, the physical radius of the apparent horizon can tend to a constant value for large cosmological time. This metric does not suffer from issues like a singular horizon for asymptotically vanishing Hubble parameter.
@@ 1 Introduction
Black holes are an indispensable part of any textbook on general relativity, which has established the fundamental mathematical tools in our understanding of the universe as a whole or in other words cosmology, and yet a consistent black hole solution (with or without a static horizon) that can be embedded in cosmology is still a matter of debate (see for instance McVittie ; Einstein:1945id ; Kaloper ; Davidson ; Faraoni ; Gaur ; Dahal ; Ion ; Poplawski ; Farrah ). Indeed there are arguments that forbid the existence of black holes with a static horizon, independent of the underlying gravitational theory Davidson Faraoni . Our main initiative in this work is to propose a way out of these no-go results and show that it is possible to find such metrics that behave like a black hole close to the horizon and tend to standard cosmology away from it, meanwhile keeping the horizon static. To this end we use the Painlevé-Gullstrand coordinates for both the black hole and cosmological metrics. In order to achieve a static horizon, we drop one of the assumptions that lead to McVittie’s metric. We let $p_{\perp}$ to be different from $p_{r}$ for radial distance $r$ small compared to the cosmological horizon (i.e. in the vicinity of the black hole). Assuming our gravitational theory is general relativity, energy density everywhere is the same as the cosmological value, and the other components also tend to cosmological values for $\frac{l_{1}}{r}\ll 1$, where $l_{1}$ is the static black hole horizon radius. This metric satisfies the null convergence condition, i.e. null energy condition (NEC) outside the horizon, but it violates it inside the black hole horizon. In order to resolve this issue we drop one of our assumptions, namely we allow the energy density to be different from the FLRW one and show that it is in principle possible to have a metric that satisfies NEC inside the black hole horizon as well.
@@
In the second part of this work we consider a more realistic setup with a constant equation of state. Specializing to the case of black hole$+$cosmic matter we find exact black hole solutions and show that it is possible to have a cosmological black hole metric which has an asymptotically (in $t$) constant physical radius. We repeat this calculation for the case of black hole$+$matter$+$cosmological constant and find similar results. In particular we show that the apparent horizons in the latter case tend to the static horizons of Schwarzschild-de Sitter metric for large $t$.
In section II, we provide an intuitive derivation of a black hole metric with a static horizon by cautiously extending the Schwarzschild-de Sitter metric to the case of cosmology with varying Hubble parameter $H$ and show that this metric does not suffer from problems like singularities at the horizon. We show that this metric could be derived via a systematic calculation with some assumptions. We consider the energy conditions for this metric assuming general relativity and find that NEC is violated inside the horizon for this proposal. We then propose a way to respect NEC inside the horizon. In section III, we consider a more realistic framework. We study the case of constant equation of state and the cases of matter and matter$+$cosmological constant separately and find that it is possible to have black hole metrics with apparent horizons that tend to a constant physical radius for large cosmological time. We conclude in section IV.
@@ 2 The feasibility of a static horizon
In this section we show that it is in principle possible to have a static black hole horizon in cosmology.
2.1 Extending the Schwarzschild-de Sitter metric
@@
where $a(t)$ is the scale factor.
This coordinate system is specially suitable for writing a metric that interpolates between the cosmological metric and the Schwarzschild metric due to the time-independent radial factor of angular coordinates.
@@
Using this analogy we propose a metric for a spherically symmetric black hole with a static horizon in cosmology as
@@
This metric has the property that the black hole and cosmological horizons are detached and it has an static horizon at $r=l_{1}$. Also if we replace $\frac{a}{\dot{a}}$ with de Sitter length $l$ for the cosmological constant, we recover the Schwarzschild-de Sitter metric.
@@
We can see that the Ricci scalar and curvature invariants like $R^{\mu\nu}R_{\mu\nu}$ and $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ remain finite at the black hole horizon. In particular we have at the horizon $r=l_{1}$
@@
Figure 1: The Ricci scalar, $R^{\mu\nu}R_{\mu\nu}$, $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ as a function of $r$, for $l_{1}=1$ and $t=10$, with blue and orange curves related to radiation and matter respectively. In the last plot, radiation and matter curves are almost coincident. These plots represent the observation that this metric has a singularity only at $r=0$ and not at the black hole horizon.
Another interesting property is that $T^{t}_{r}$ which we find from requiring Einstein’s equation is zero. In other words there is no flux of energy in radial direction. This is the requirement also present in McVittie’s solution. On the other hand, the large $r$ limit of stress tensor tends to the cosmological values.
@@
2.2 A systematic derivation
@@
$b)$ There is a static horizon at $r=l_{1}$, such that $g_{tt}|_{r=l_{1}}=0$.
@@
We note that compared to McVittie’s assumptions, we have dropped the $p_{\perp}=p_{r}$ assumption and instead we have added assumption b.
@@ 2.3 Energy conditions
For a typical power-law scale factor, this is always positive outside the horizon and negative inside the horizon (since $g_{tr}<0$). Therefore assuming the gravitational theory is general relativity, the stress tensor corresponding to this metric does not satisfy the weak and null energy conditions inside the black hole horizon.
Considering the strong energy condition we find that it is violated in the vicinity of the black hole (see figure 2).
Figure 2: Here we have plotted $T_{tt}-\frac{1}{2}g_{tt}T$ as a function of $r$. We have chosen $l_{1}=1$ and $t=10$. The blue and orange curves respectively correspond to radiation and matter dominated scale factors. The strong energy condition is violated in the black hole’s vicinity.
@@ 3 The case of constant equation of state
In this section we consider the possibility that the black hole metric respects the cosmological stress tensor equation of state. Assuming $T^{t}_{r}=0$, from (19) we find
@@
On the other hand from $G^{t}_{t}=-\kappa_{4}\rho$ and (35) we find
@@
which is the Schwarzschild metric in Painlevé-Gullstrand coordinates (This is evident by changing $t$ to $t^{\prime}$ such that $dt^{\prime}=\sqrt{\gamma(t)}dt$). Therefore, in order to find a cosmological black hole metric, our preference will be $\gamma(t)=1$ in what follows.
@@ 3.1 The case of $\omega\neq 0$
Subtracting (40) from (42) we further find
@@
$03\zeta^{2}(t)r^{2}+4r\zeta(t)\alpha(r)+2r^{2}\zeta(t)\alpha^{\prime}(r)+\frac{2}{1+\omega}r^{2}\dot{\zeta}+\alpha^{2}(r)\\ +2r\alpha(r)\alpha^{\prime}(r)=$ 0 . (44)
The only nontrivial solution of this equation which is of cosmological interest is $\alpha(r)=0$, $\zeta(t)=\frac{2}{3(1+\omega)t}$. Therefore we do not have a cosmological black hole solution that satisfies $p_{r}=p_{\perp}=\omega\rho$ for $\omega\neq 0$.
3.2 The case of matter with $\omega=0$
@@
In this section we consider the case of a Schwarzschild black hole immersed in cosmological matter. In this case we have $p_{r}=p_{\perp}=0$. From stress energy conservation we can find the structure of $\rho$. We have
@@
and this also satisfies (48). Using (51) this means
@@
where in the last equality we used $r=R\,a$. We see that this simple case corresponds to the Schwarzschild metric in Painlevé-Gullstrand coordinates. We also note that in this coordinate the black hole singularity is at the surface $\frac{3}{2}t-b(R)=0$. There is a freedom in our choice of $b(R)$.
Now returning to the case of black hole in matter-dominated cosmology we can add a term related to the cosmological matter density to $\beta(R)$. We note that in the asymptotic case we should have $\kappa_{4}\beta(R)=\beta_{0}$ where $\beta_{0}$ is a constant. Considering that this is also the case for large $t$ we set $\kappa_{4}\beta(R)=\beta_{0}$ and find
@@
As an example, if we choose $b(R)=\frac{\kappa l_{s}^{4}}{R^{3}}-d$, using (59) and (61) we can find
@@
For large $t$ one of these tends to $r_{H}=\frac{3t}{2}$, which is the cosmological horizon, and the other tends to $r_{H}=l_{s}$, which is the black hole horizon. Therefore, for large $t$ we effectively have a static black hole horizon.
We had freedom in our choice of $b(R)$. Requiring that the metric tends to the cosmological metric for large $R$ will force us to choose $b(R)$ such that it tends to a constant for large $R$. The surface $\frac{3t}{2}=b(R)$, $r=0$ is where both the cosmological singularity and black hole singularity reside. We see that the choice for $b(R)$ fixes both the singularity and the evolution of the black hole in $t$.
We can prove that generically the resulting apparent horizon in this case will be dynamical (apart from the asymptotic behavior). To see this we write
@@
From this, we can see that it is impossible to have $\dot{r}_{H}=0$ exactly. [^2] To have $\dot{r}_{H}=0$ from the first equality we should either have $\dot{R}_{H}=0$ which contradicts with the second equality or we should have $\beta(R_{H})=0$ which again leads to $R_{H}=const.$ which is a contradiction.
3.3 The case of matter$+$cosmological constant
@@
The smallest root of this equation is the black hole horizon. This means that the location of the black hole horizon for large $t$ tends to its Schwarzschild-de Sitter counterpart.

### v2

@@ Abstract
Although previous results have ruled out the possibility of a static horizon in cosmology, we present black hole and white hole metrics that retain static horizons while reproducing cosmological behavior at large distances. Using an appropriate coordinate choice, we demonstrate that a static horizon can exist in a cosmological setting without introducing curvature invariant singularities at the horizon. The resulting metric reduces to the Schwarzschild–de Sitter solution when the Hubble parameter is constant. We find that white hole metrics in an expanding universe (or black holes in a contracting universe) are significantly easier to construct, as a black hole in an expanding cosmology requires the velocity function to change sign. Consequently, this work primarily examines white holes in expanding cosmologies as a foundation for subsequent analysis of black holes in expanding universes. In later sections, we investigate scenarios involving a white hole coupled with cosmological matter, as well as a white hole with both matter and a cosmological constant. Assuming the pressure component takes its cosmological value, we show that the physical radius of the apparent horizon can asymptotically approach a constant value at late times. This metric avoids pathologies such as a singular horizon in the limit of a vanishing Hubble parameter. Finally, we analyze the realistic case of a black hole embedded in pressure-less cosmological matter with and without a cosmological constant and explore its properties.
@@ 1 Introduction
Black holes are an indispensable part of any textbook on general relativity, which has established the fundamental mathematical tools in our understanding of the universe as a whole or in other words cosmology, and yet a consistent black hole solution (with or without a static horizon) that can be embedded in cosmology is still a matter of debate (see for instance McVittie ; Einstein:1945id ; Kaloper ; Davidson ; Faraoni ; Gaur ; Dahal ; Ion ; Poplawski ; Farrah ; Cadoni1 ; Cadoni2 ). Indeed there are arguments that forbid the existence of black holes with static horizons, independent of the underlying gravitational theory Davidson Faraoni . Our main initiative in this work is to propose a way out of these no-go results and show that it is possible to find such metrics that behave like a black hole close to the horizon and tend to standard cosmology away from it, meanwhile keeping the horizon static. To this end, in section II, we use the Painlevé-Gullstrand coordinates for both the black hole and cosmological metrics. Due to the simplicity of deriving a white hole in expanding cosmology (or similarly, a black hole in contracting cosmology), we first consider this case and study the possibility of a static horizon. In order to achieve a static horizon, we drop one of the assumptions that led to McVittie's metric. We let $p_{\perp}$ be different from $p_{r}$ for radial distance $r$ small compared to the cosmological horizon (i.e., in the vicinity of the white hole). Assuming our gravitational theory is general relativity, energy density everywhere is the same as the cosmological value, and the other components also tend to cosmological values for $\frac{l_{1}}{r}\ll 1$, where $l_{1}$ is the static white hole horizon radius. This metric satisfies the null convergence condition, i.e. null energy condition (NEC) outside the horizon, but it violates it inside the white hole horizon. To resolve this issue, we drop another one of our assumptions, namely, we allow the energy density to be different from the FLRW one, and show that it is, in principle, possible to have a metric that satisfies NEC inside the white hole horizon as well. We then consider the case of a black hole in expanding cosmology that requires a sign change in the velocity function $h$ by considering a more generic metric ansatz.
@@
In the second part of this work (section III), we consider a more realistic setup with a constant equation of state. Specializing to the case of white hole plus cosmic matter in an expanding cosmology, we find exact white hole solutions and show that it is possible to have a cosmological white hole metric which has an asymptotically (in $t$) constant physical radius. We repeat this calculation for the case of white hole plus matter, plus cosmological constant, and find similar results. In particular, we show that the apparent horizons in the latter case tend to the static horizons of the Schwarzschild-de Sitter metric for large $t$. We then consider a more generic metric ansatz that includes a curvature-like behavior (relative to a flat FLRW metric) and find a black hole metric in matter-dominated expanding cosmology that has a sign change in the velocity function $h$. We conclude in Section IV.
@@ 2 The feasibility of a static horizon
In this section, we show that it is in principle possible to have a static horizon in cosmology.
2.1 A naive extension of the Schwarzschild-de Sitter metric
@@
where $a(t)$ is the scale factor. This coordinate system is specially suitable for writing a metric that interpolates between the cosmological metric and the Schwarzschild metric due to the time-independent radial factor of angular coordinates.
@@
We should note that due to a negative $g_{tr}$ component, this metric in fact represents a white hole in these coordinates Volovik . Using this analogy we propose a metric for a spherically symmetric white hole with a static horizon in cosmology as
@@
This metric has the property that the white hole and cosmological horizons are detached and it has an static horizon at $r=l_{1}$. Also if we replace $\frac{a}{\dot{a}}$ with de Sitter length $l$ for the cosmological constant, we recover the Schwarzschild-de Sitter metric.
@@
We can see that the Ricci scalar and curvature invariants like $R^{\mu\nu}R_{\mu\nu}$ and $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ remain finite at the horizon. In particular we have at the horizon $r=l_{1}$
@@
Figure 1: The Ricci scalar, $R^{\mu\nu}R_{\mu\nu}$, $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ as a function of $r$, for $l_{1}=1$ and $t=10$, with blue and orange curves related to radiation and matter respectively. In the last plot, radiation and matter curves are almost coincident. These plots represent the observation that this metric has a singularity only at $r=0$ and not at the event horizon.
Another interesting property is that $T^{t}_{r}$ which we find from requiring Einstein's equation is zero. In other words there is no flux of energy in radial direction. This is the requirement also present in McVittie's solution. On the other hand, the large $r$ limit of stress tensor tends to the cosmological values.
@@
2.2 A systematic derivation for the white hole case
@@
$b)$ There is a static horizon at $r=l_{1}$, such that $g^{rr}|_{r=l_{1}}=0$.
@@
We note that compared to McVittie's assumptions, we have dropped the $p_{\perp}=p_{r}$ assumption and instead we have added assumption b.
@@ 2.3 Energy conditions
For a typical power-law scale factor, this is always positive outside the horizon and negative inside the horizon (since $g_{tr}<0$). Therefore assuming the gravitational theory is general relativity, the stress tensor corresponding to this metric does not satisfy the weak and null energy conditions inside the event horizon.
Considering the strong energy condition we find that it is violated in the vicinity of the white hole (see figure 2).
Figure 2: Here we have plotted $T_{tt}-\frac{1}{2}g_{tt}T$ as a function of $r$. We have chosen $l_{1}=1$ and $t=10$. The blue and orange curves respectively correspond to radiation and matter dominated scale factors. The strong energy condition is violated in the white hole's vicinity.
@@ 2.4 NEC and static horizon
2.5 The feasibility of a static horizon, the black hole case
In this section, we consider a more generic form of the metric that can represent a black hole in cosmology with a static horizon. For this, we consider the metric ansatz Volovik
$ds^{2}=-\big{(}N^{2}-g_{rr}h^{2}\big{)}dt^{2}+g_{rr}dr^{2}+2g_{rr}h\,dtdr+r^{2}d\Omega_{2}^{2},$ (33)
We assume $N^{2}=\frac{1}{g_{rr}}=\lambda(t,r)$.
Considering the assumption that $G^{t}_{t}=-3H^{2}$, which is its cosmological value, we can write
$\partial_{r}(\frac{rh^{2}}{\lambda}+r-r\lambda)=H^{2}\partial_{r}(r^{3}),$ (34)
$h^{2}=\lambda(\lambda-1+H^{2}r^{2}+\frac{\alpha(t)}{r})$ (35)
Using the requirement that we have a static horizon at $r=l_{1}$, we find
$g^{rr}=\lambda(r)-\frac{h^{2}}{\lambda}|_{r=l_{1}}=0.$ (36)
From this we have $\alpha(t)=l_{1}(1-H^{2}l_{1}^{2})$ and
$h^{2}=\lambda\big{(}\lambda-1+H^{2}r^{2}+\frac{l_{1}}{r}(1-H^{2}l_{1}^{2})\big{)}.$ (37)
In accordance with Volovik , we can choose
$\lambda=1-3(\frac{l_{s}H}{2})^{2/3},\qquad l_{s}=l_{1}(1-H^{2}l_{1}^{2}).$ (38)
Then we have
$h^{2}=(1-\lambda)\lambda(r-r_{0})^{2}\frac{r+2r_{0}}{3\,r\,r_{0}^{2}},\qquad r_{0}=(\frac{l_{s}}{2H^{2}})^{1/3}.$ (39)
Therefore we can choose different signs of the square root for $r>r_{0}$ and $r<r_{0}$.
Again one can show that this result violates NEC. Similar to the white hole case we can drop the assumption that the energy density is equal to the cosmological value and find a metric that also satisfies NEC.
@@ 3 The case of constant equation of state
In this section we consider the possibility that the metric respects the cosmological stress tensor equation of state.
3.1 The white hole case
Here we use the metric ansatz (18). Assuming $T^{t}_{r}=0$, from (19) we find
@@ 3.1 The white hole case
On the other hand from $G^{t}_{t}=-\kappa_{4}\rho$ and (42), we find
@@
which is the Schwarzschild white hole metric in Painlevé-Gullstrand coordinates (This is evident by changing $t$ to $t^{\prime}$ such that $dt^{\prime}=\sqrt{\gamma(t)}dt$). Therefore, in order to find a cosmological white hole metric, our preference will be $\gamma(t)=1$ in what follows.
@@ 3.1.1 The case of $\omega\neq 0$
Subtracting (47) from (49), we further find
@@
$3\zeta^{2}(t)r^{2}+4r\zeta(t)\alpha(r)+2r^{2}\zeta(t)\alpha^{\prime}(r)+\frac{2}{1+\omega}r^{2}\dot{\zeta}+\alpha^{2}(r)\\ +2r\alpha(r)\alpha^{\prime}(r)=0$ (51)
The only nontrivial solution of this equation which is of cosmological interest is $\alpha(r)=0$, $\zeta(t)=\frac{2}{3(1+\omega)t}$. Therefore we do not have a cosmological solution that satisfies $p_{r}=p_{\perp}=\omega\rho$ for $\omega\neq 0$.
3.1.2 The case of matter with $\omega=0$
@@
In this section we consider the case of a Schwarzschild white hole immersed in cosmological matter. In this case we have $p_{r}=p_{\perp}=0$. From stress energy conservation we can find the structure of $\rho$. We have
@@
which also satisfies (55). Using (58) this means
@@
where in the last equality we used $r=R\,a$. We see that this simple case corresponds to the Schwarzschild metric in Painlevé-Gullstrand coordinates. We also note that in this coordinate the singularity is at the surface $\frac{3}{2}t-b(R)=0$. There is a freedom in our choice of $b(R)$.
Now returning to the case of matter-dominated cosmology we can add a term related to the cosmological matter density to $\beta(R)$. We note that in the asymptotic case we should have $\kappa_{4}\beta(R)=\beta_{0}$ where $\beta_{0}$ is a constant. Considering that this is also the case for large $t$ we set $\kappa_{4}\beta(R)=\beta_{0}$ and find
@@
As an example, if we choose $b(R)=\frac{\kappa l_{s}^{4}}{R^{3}}-d$, using (66) and (68) we find
@@
For large $t$ one of these tends to $r_{H}=\frac{3t}{2}$, which is the cosmological horizon, and the other tends to $r_{H}=l_{s}$, which is the white hole horizon. Therefore, for large $t$ we effectively have a static white hole horizon.
We had freedom in our choice of $b(R)$. Requiring that the metric tends to the cosmological metric for large $R$ will force us to choose $b(R)$ such that it tends to a constant for large $R$. The surface $\frac{3t}{2}=b(R)$, $r=0$ is where both the cosmological singularity and white hole singularity reside.
We can prove that generically the resulting apparent horizon in this case will be dynamical (apart from the asymptotic behavior). To see this, we write
@@
From this, we can see that it is impossible to have $\dot{r}_{H}=0$ exactly. [^2] To have $\dot{r}_{H}=0$ from the first equality we should either have $\dot{R}_{H}=0$ which contradicts the second equality or we should have $\beta(R_{H})=0$ which again leads to $R_{H}=const.$ which is a contradiction.
3.1.3 The case of matter plus cosmological constant
@@
The smallest root of this equation is the white hole horizon. This means that the location of the white hole horizon for large $t$ tends to its Schwarzschild-de Sitter counterpart.
@@
3.2 The Black Hole case
In this section we study the case of black hole immersed in cosmological matter and matter plus cosmological constant.
3.2.1 The case of matter
Beginning from the metric ansatz
$ds^{2}=-N^{2}dt^{2}+g_{rr}(dr-hdt)^{2}+r^{2}d\Omega_{2}^{2},$ (85)
we find after a change of variable $r=f(t,R)$ such that $h=\dot{f}$ the metric
$ds^{2}=-N^{2}dt^{2}+g_{rr}(f^{\prime})^{2}dR^{2}+f^{2}d\Omega_{2}^{2}.$ (86)
Assuming we only have a dust contribution to stress tensor such that only $T^{t}_{\,t}=-\kappa_{4}\rho$ is nonzero, we find from stress tensor conservation
$N^{\prime}=0.$ (87)
Therefore $N$ is only a function of $t$ and by redefining $t$ we can set it to $N=1$.
Requiring $T^{t}_{\,R}=0$ we further find $\dot{g}_{rr}=0$ and we define $g_{rr}=\frac{1}{\lambda^{2}(R)}$. From $T^{R}_{\,R}=0$ we find
$\partial_{t}\big{(}(1-\lambda^{2})f+f\dot{f}^{2}\big{)}=0,$ (88)
On the other hand from conservation equation we can fix the form of $\kappa_{4}\rho$ to be
$\kappa_{4}\rho=\frac{\beta(R)}{f^{2}f^{\prime}}$ (89)
and from $T^{t}_{\,t}=-\kappa_{4}\rho$ we find
$\partial_{R}\big{(}(1-\lambda^{2})f+f\dot{f}^{2}\big{)}=\beta(R).$ (90)
Therefore we can write this as
$f\dot{f}^{2}=\zeta(R)-\omega(R)f$ (91)
where we have defined $\zeta(R)$ such that $\zeta^{\prime}(R)=\beta(R)$ and $\omega(R)=1-\lambda^{2}(R)$.
The general solution to this equation can be written as
$\frac{\zeta}{\omega\sqrt{\omega}}\arcsin{\sqrt{\frac{\omega f}{\zeta}}}-\frac{\sqrt{f(\zeta-\omega f)}}{\omega}=\pm t+b_{\pm}(R).$ (92)
We note that (91) requires $f<\frac{\zeta}{\omega}$ for $\omega>0$.
We study the behavior of this solution in different limits. For $f\ll\frac{\zeta}{\omega}$ the solution becomes
$f=\zeta^{1/3}(R)(\frac{3}{2}(\pm t+b_{\pm}(R)))^{2/3}.$ (93)
In the limit where $f\approx\frac{\zeta}{\omega}$ the first trem in (92) is dominant and the solution takes the form
$f=\frac{\zeta}{\omega}\sin^{2}\big{(}\frac{\omega^{3/2}}{\zeta}(\pm t+b_{\pm}(R))\big{)},$ (94)
such that at ±$\pm t+b_{\pm}(R)=\frac{\pi\,\zeta(R)}{2\,\omega^{3/2}(R)}$, $f$ reaches its maximum for a given $R$.
To understand the behavior of the solution we define $f=\frac{\zeta^{1/3}}{d}\hat{f}$ where $d$ has dimension length and assume $\omega(R)=(\frac{\zeta}{d})^{2/3}$ to find
$\hat{f}\dot{\hat{f}}^{2}=d-\hat{f},$ (95)
with solutions
$d\arcsin{\sqrt{\frac{\hat{f}}{d}}}-\sqrt{\hat{f}(d-\hat{f})}=\pm t+b_{\pm}(R).$ (96)
For a fixed $R$, $b_{\pm}(R)$ are constants. If we set $b_{\pm}=0$, then the solutions are as in figure 3.
[…]
To have a reasonable solution that moves from one solution smoothly to the other, we note that $\hat{f}$ takes its maximum value at ±$\pm t^{c}_{\pm}+b_{\pm}=\frac{\pi\,d}{2}$ where $t^{c}_{\pm}$ are transition times. Therefore for a smooth transition we need to have $t^{c}_{-}=t^{c}_{+}=t^{c}$ from which we find $b_{+}+b_{-}=\pi\,d$. For $b_{+}=0$ and $b_{-}=\pi\,d$ the plot becomes as in figure 4.
Figure 4: The plot for $\hat{f}$ with $b_{\pm}$ chosen such that we have a smooth transition between the $+$ and -- branches.
According to this, the behavior of $\hat{f}$ is such that at each value of $R$, for early time ($t<t^{c}=\frac{\pi\,\zeta}{2\,\omega^{3/2}}$) we have an increasing scale factor reminiscent of a an expanding cosmology. After the crossover time the scale factor becomes decreasing which is reminiscent of the behavior in the vicinity of a black hole.
According to (93) the metric away from the transition point only depends on $\zeta$ and $b_{\pm}$. Therefore our calculations regarding the apparent horizon remain the same as the white-hole case and we can have an apparent horizon which tends to a constant for large time.
3.2.2 The case of matter plus cosmological constant
Similar to the previous section we can again set $N=1$ and from $T^{t}_{R}=0$ we find $\dot{\lambda}=0$. From $\kappa_{4}T^{R}_{R}=-\frac{3}{l_{0}^{2}}$ related to the cosmological constant we can write
$\partial_{t}\big{(}(1-\lambda^{2}(R))f+f\dot{f}^{2}-\frac{f^{3}}{l_{0}^{2}}\big{)}=0.$ (97)
On the other hand from energy momentum conservation we find as before
$\kappa_{4}\rho=\frac{\beta(R)}{f^{2}f^{\prime}}.$ (98)
From $\kappa_{4}T^{t}_{t}=-\kappa_{4}\rho-\frac{3}{l_{0}^{2}}$ we find
$(1-\lambda^{2})f+f\dot{f}^{2}-\frac{f^{3}}{l_{0}^{2}}=\zeta(R),$ (99)
where we have $\zeta^{\prime}(R)=\beta(R)$.
We study this equation in two limits. Close to the black hole we expect from previous results that $f\ll\frac{\zeta}{\omega}$ (defining $\omega=1-\lambda^{2}$). In this limit we have
$f\dot{f}^{2}-\frac{f^{3}}{l_{0}^{2}}\approx\zeta(R),$ (100)
and we find
$f=(l_{0}^{2}\zeta)^{1/3}\sinh^{2/3}\big{(}\pm\frac{3t}{2l_{0}}+\frac{b(R)}{l_{0}}\big{)}.$ (101)
Another important limit is where $f\approx\frac{\zeta}{\omega}$. This is close to the region where the sign of $h=\dot{f}$ changes. In order to understand the behavior of $f$ in this region, we find it suitable to set define $f=\big{(}\frac{\zeta}{d}\big{)}^{1/3}\hat{f}$ and choose $\omega=\big{(}\frac{\zeta}{d}\big{)}^{2/3}$ to find
$\hat{f}+\hat{f}\dot{\hat{f}}^{2}-\frac{\hat{f}^{3}}{l_{0}^{2}}=d.$ (102)
To study the region close to $f=\frac{\zeta}{\omega}$ we write $\hat{f}=d+\phi$ where $\phi\ll d$ and find (we assume $d\ll l_{0}$)
$\phi+d\,\dot{\phi}^{2}\approx\frac{d^{3}}{l_{0}^{2}},$ (103)
with solution
$\phi=\frac{d^{3}}{l_{0}^{2}}-\frac{d}{4}\big{(}\frac{t}{d}-l_{0}c(R)\big{)}^{2}.$ (104)
Here $c(R)$ is an arbitrary function of $R$. We see that this has the same behavior as in figure 4 close to the cross over point $t_{c}=l_{0}\,d\,c(R)$.